How much heat is released from a 500 g iron rod when it cools from 125°C to 20°C? The specific heat of Fe is 0.451 J/gºC
Answer #1

\bf Given:

\bf m_{Fe}=500\;g, T_1=125^oC,T_2=20^oC,C_{Fe}=0.451\;J/g^oC

Heat released can be calculated by the formula

\bf Q=mC\Delta T

\bf Q= m_{Fe}\times C_{Fe} \times (T_2-T_1)

\bf Q= 500\;g\times 0.451\;J/g^oC \times (20^oC-125^oC)

\bf Q= 500\;g\times 0.451\;J/g^oC \times (-105^oC)

\LARGE \bf Q= -23677.5\;J

\LARGE \bf negative\text{ sign indicates that }Heat\;is\;lost